# arXiv:2602.20833v2[cs.DS] 26 Feb 2026

### v2

@@ 8 Conclusion
We presented the DRESS family, a hierarchy of continuous structural refinement algorithms. Starting from Original-DRESS—a single equation proposed in 2018—we showed it distinguishes graphs that 1-WL cannot (Theorem 1). We then generalized it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions. Generalized-DRESS abstracts further, opening the aggregation function and norm as free parameters. $\Delta$-DRESS extends the framework by running DRESS on each node-deleted subgraph, connecting it to the Kelly–Ulam reconstruction conjecture and empirically distinguishing Strongly Regular Graphs that confound 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs. These results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL, without the prohibitive computational cost of higher-order WL tests. An open-source implementation with bindings for C, C++, Python, Rust, Go, Julia, R, MATLAB, and WebAssembly is publicly available at https://github.com/velicast/dress-graph.

### v3

@@ 8 Conclusion
We presented the DRESS family, a hierarchy of continuous structural refinement algorithms. Starting from Original-DRESS–a single equation proposed in 2018–we showed it distinguishes graphs that 1-WL cannot (Theorem 1). We then generalized it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions. Generalized-DRESS abstracts further, opening the aggregation function and norm as free parameters. $\Delta$-DRESS extends the framework by running DRESS on each vertex-deleted subgraph, connecting it to the Kelly–Ulam reconstruction conjecture, and empirically distinguishes Strongly Regular Graphs that confound 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs. These results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL, without the prohibitive computational cost of higher-order WL tests. An open-source implementation with bindings for C, C++, Python, Rust, Go, Julia, R, MATLAB, and WebAssembly is publicly available at https://github.com/velicast/dress-graph.